4. What is the molar mass of a compound if a gaseous sample has a density of 1.173 g/L at 20 °C and 630 torr? The temperature in Celsius is known to two significant figures.

5. A 1.50-L bulb containing Ne at 125 torr is connected by a valve to a 2.50-L bulb containing CH4 at 650 torr. The valve between the two bulbs is opened and the two gases mix. The initial gas pressures as known to three significant figures.

(a) What is the partial pressure (torr) of Ne? (b) What is the partial pressure (torr) of CH4? (c) What is the total pressure?
(d) What is the mole fraction of Ne?

  1. Consider the following: SiH3Cl, O2, CF3H, and CF4.
    (a) Which one will diffuse the fastest?
    (b) What is the rms speed (m/s) of the slowest-moving molecule at 25°C?
Answer #1

4)

P= 630.0 torr

= (630.0/760) atm

= 0.8289 atm

T= 20.0 oC

= (20.0+273) K

= 293 K

density = 1.173 g/L

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

0.8289 atm * MM = 1.173g/L * 0.08206 atm.L/mol.K *293.0 K

MM = 34.0228 g/mol

Answer: 34.0 g/mol

Only 1 question at a time please